# Approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search

### v1

@@ 4 Convergence Analysis
Under Assumption 4.1, since $\tilde{\mathcal{T}}_{\lambda}(x)$ is strongly convex and closed, it has a unique minimizer.
@@ 4.1 Properties of Proposed Algorithm
Therefore, substituting $y\leftarrow x+d$ on the above inequality, we obtain
@@ 4.2 Global Subsequential Convergence
Assumption 4.5(i) is often assumed in nonsmooth optimization when the problem includes nonsmooth lower semicontinuous functions [Bolte-2018, Takahashi]. In fact, a lower semicontinuous, level-bounded, and proper function has a minimum [rockafellar2009variational, Theorem 1.9]. Assumption 4.5(ii) would often hold when the influence of $f$ is dominant compared to that of $g$. We will discuss this issue for more details in Section A. Moreover, under Assumption 4.5(ii), Assumption 4.5(iii) always holds because $\Psi$ is bounded below from Assumption 3.1(iv) and the right-hand side of (3.6) is unbounded below. In this case, we have $\|t_{k}d^{k}\|\to 0$.
@@
where the last inequality holds because $\phi$ is $\sigma$-strongly convex. Let $\bar{x}\in\mathbb{R}^{n}$ be an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$ and let $\{x^{k_{j}}\}_{j\in\mathbb{N}}$ be a subsequence such that $x^{k_{j}}\to\bar{x}$ by Bolzano–Weierstrass theorem. From (4.26) and Cauchy–Schwartz inequality, we have
@@
If $\|\nabla f(x^{k})+\xi^{k}\|=0$, then $x^{k}$ becomes a stationary point. We assume $\|\nabla f(x^{k})+\xi^{k}\|>0$. By the triangle inequality, we have $\|\nabla f(x^{k})+\xi^{k}\|\leq\|\nabla f(x^{k})\|+\|\xi^{k}\|$. Since the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded, $\|\nabla f(x^{k})\|$ is bounded by the extreme value theorem and $\|\xi^{k}\|$ is also bounded (see, e.g.[Takahashi2022-ml, Theorem 1(ii)]). Thus, $\|\nabla f(x^{k})+\xi^{k}\|$ is bounded, i.e., due to (4.27), $d^{k}$ is also bounded. Thus, there exists a subsequence $\{d^{k_{j}}\}_{j\in\mathbb{N}}$ such that $d^{k_{j}}\to\bar{d}$ as $j\to\infty$ by Bolzano–Weierstrass theorem. Then, by Lemma 4.6, the sequence $\{x^{k_{j}}+t_{k_{j}}d^{k_{j}}\}_{j\in\mathbb{N}}$ also converges to $\bar{x}$.
@@
which induces $d^{k_{j}}\to 0$. Because $f$ and $g$ are lower semicontinuous, from (4.24), we have
@@
where the first inequality holds due to Cauchy–Schwartz inequality and $\nabla f$ being Lipschitz continuous on any compact subset from Assumption 4.8, and the second inequality holds due to (4.34). Moreover, using the inequality (4.26), we have
@@ 4.3 Global Convergence
where the second inequality holds because of the Lipchitz continuity of $\nabla f$ and $\nabla\phi$ on compact subsets from Assumption 4.8 and Assumption 3.1(i), and the last inequality holds from line 1 in Algorithm 1.
@@
    For any $k\geq 1$ and some $\bar{k}_{0}\in\mathbb{N}$, the following conditions hold:
@@
Suppose that Assumptions 3.1, 3.2, 4.1, 4.5, and 4.8 hold. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by ABPG-VMAW and let $\bar{x}$ be a stationary point of (1.1). Suppose further that $\Psi$ is a KL function with $\psi$ in the KL inequality (2.5) taking form $\psi(s)=cs^{1-\theta}$ for some $\theta\in[0,1)$ and $c>0$. Then, the following statements hold:
@@ 5 Numerical Experiments
5.1 $\ell_{p}$ Regularized Least Squares Problem
We consider the sparse $\ell_{p}$ (whose $p$ is slightly larger than 1) regularized least squares problem [doi:10.1137/18M1194456, 7472557]:
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}\|Ax-b\|^{2}+\frac{\theta_{p}}{p}\|x\|^{p}_{p},$ (5.1)
where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $\theta_{p}>0$. Let $g\equiv 0$. We also use $f$ and $\phi$ give by
@@ 5.1 $\ell_{p}$ Regularized Least Squares Problem
Note that $f$ and $\phi$ are $\mathcal{C}^{1}$ if $p>1$ while $\nabla f$ and $\nabla\phi$ are not globally Lipschitz continuous. Although we can choose any $\lambda>0$, we use $\lambda$ given by $\lambda<1/L$ if $(f,\phi)$ is $L$-samd (see, for more details, Remark 3.4). Note that our algorithm does not require the $L$-smad property.
@@
We compare ABPG-VMAW with ABPG [Takahashi], the proximal gradient algorithm (PG) with a constant step-size, and PG with line search (PGL). We set $c_{1}=0.99$, $c_{2}=0.999$, $\mu=0.9$, and $\eta=2$ for ABPG-VMAW and $c_{1}=0.99$ and $\delta=0.9$ for ABPG. Although $\nabla f$ is not Lipschitz continuous, PG uses the step-size $1/L$ given by (5.3). Note that PG does not guarantee global convergence. PGL searches $\lambda_{k}>0$ satisfying the descent lemma and uses the initial step-size $\lambda_{0}=1/L$ given by (5.3[doi:10.1137/1.9781611974997, p.283]. The initial point $x^{0}\in\mathbb{R}^{n}$ is generated from i.i.d. normal distribution. The maximum iteration is 1000. The terminal condition is $\|x^{k}-x^{k-1}\|\leq 10^{-8}$.
@@
Next, we show the average performance of the four methods—ABPG-VMAW, ABPG, PG, and PGL—on the $\ell_{p}$-regularized least squares problem. Specifically, we selected combinations of $m$ and $n$ from the sets $\{100,200\}\times\{1000,2000,5000\}$, and for each combination, we generated a random $m\times n$ matrix, the ground truth $x^{*}\in\mathbb{R}^{n}$, which has $10\%$ nonzero elements, from i.i.d. normal distribution $100$ times. For each generated instance, we set $b=Ax^{*}$, $p=1.2$, and $\theta_{p}=0.1$. Table 1 presents the average performance, including the number of iterations, the accuracy of the recovered point, the objective values, and computation time, across 100 different instances. ABPG-VMAW outperformed ABPG, PG, and PGL. Moreover, ABPG-VMAW converged in fewer iterations and in a shorter amount of time than ABPG, PG, and PGL.
Table 1: Average number of iterations, objective function value, and accuracy, CPU time for ABPG-VMAW, ABPG, PG, and PGL using 100100 random instances of $\ell_{p}$-regularized least squares problem (5.1)
@@
100 1000 ABPG-VMAW 91 0.489 0.990 0.068
ABPG 804 0.489 0.990 0.524
PG 1000 5.996 1.092 0.179
PGL 1000 11.054 3.289 0.220
2000 ABPG-VMAW 155 0.498 0.995 0.387
ABPG 894 0.498 0.995 1.322
PG 1000 6.371 1.055 0.849
PGL 1000 25.675 5.407 0.923
5000 ABPG-VMAW 161 0.497 0.999 2.038
ABPG 1000 0.497 0.999 6.336
PG 1000 6.377 1.025 5.489
PGL 1000 80.069 10.428 5.668
200 1000 ABPG-VMAW 54 0.492 0.994 0.052
ABPG 697 0.492 0.994 0.742
PG 1000 9.565 1.226 0.215
PGL 1000 8.930 2.651 0.276
2000 ABPG-VMAW 88 0.496 0.998 0.234
ABPG 724 0.496 0.998 1.405
PG 1000 11.200 1.167 0.860
PGL 1000 21.725 4.550 0.959
5000 ABPG-VMAW 133 0.495 0.999 1.809
ABPG 858 0.495 0.999 6.251
PG 1000 12.519 1.096 5.620
PGL 1000 68.325 9.025 5.846
@@ 5.2 Nonnegative Linear Inverse Problem
Given a nonnegative matrix $A\in\mathbb{R}^{m\times n}_{+}$ and a nonnegative vector $b\in\mathbb{R}^{m}_{+}$, the goal of nonnegative linear inverse problems is to recover a signal $x\in\mathbb{R}^{n}_{+}$ such that $Ax\simeq b$. Nonnegative linear inverse problems have been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de] as well as in optimization [Bauschke2017-hg, Takahashi]. To achieve the goal of nonnegative linear inverse problems, we focus on the convex optimization problem given by
$\min_{x\in\mathbb{R}^{n}_{+}}D_{\rm{KL}}(Ax+b)+\theta_{1}\|x\|_{1},$ (5.4)
@@
Let $f(x)=D_{\rm{KL}}(Ax,b)$ and $g(x)=\theta_{1}\|x\|_{1}$. We use $\phi_{0}(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ as a kernel generating distance for BPG and $\phi_{1}(x)=\phi_{0}+\frac{1}{2}\|x\|^{2}$ as one for our algorithm and ABPG. In this case, we also define $C=\operatorname{int}\operatorname{dom}\phi_{0}=\operatorname{int}\operatorname{dom}\phi_{1}=\mathbb{R}^{n}_{+}$. When $\sum_{i=1}^{m}a_{ij}=1$, the pair $(f,\phi_{0})$ is $1$-smad [Bauschke2017-hg] and the pair $(f,\phi_{1})$ is also $1$-smad [Takahashi]. We compare ABPG-VMAW with ABPG [Takahashi], PGL, and BPG. Those subproblems are solved in closed form.
The problem setting is as follows. We generate the matrix $A\in\mathbb{R}^{n\times m}$ and the ground truth $x^{*}\in\mathbb{R}^{n}$, which has 5% nonzero elements, from i.i.d. normal distribution. We set $b=Ax^{*}$. For $(n,m)=(200,500)$ and $\theta_{1}=0.05$, Figure 3 shows the objective function value $\Psi(x^{k})$ and the accuracy $\|x^{k}-x^{*}\|$ at each iteration on a logarithmic scale and Figure 4 shows those on the time axis.
@@
Figure 3: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (red) on the nonnegative linear inverse Problem (5.4)
@@
Figure 4: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (red) on the nonnegative linear inverse problem (5.4)

### v2

@@ 4 Convergence Analysis
Under Assumption 4.1, since $\tilde{\mathcal{T}}_{\lambda}(x)$ is strongly convex and closed, it has a unique minimizer. If $\phi$ is convex but not strongly convex, one can enforce strong convexity by adding the quadratic term $\frac{1}{2}\|\cdot\|^{2}$. When $g\equiv 0$, this modification does not affect the difficulty of the subproblems. When $g\not\equiv 0$, we are merely adding a quadratic and separable term $\frac{1}{2}\|\cdot\|^{2}$ to the subproblem that already contains a quadratic term. Therefore, the subproblems remain relatively easy to solve.
@@ 4.1 Properties of Proposed Algorithm
Therefore, substituting $y\leftarrow x+d$ into the above inequality, we obtain
@@ 4.2 Global Subsequential Convergence
Assumption 4.5(i) is often assumed in nonsmooth optimization when the problem includes nonsmooth lower semicontinuous functions [Bolte-2018, Takahashi]. In fact, a lower semicontinuous, level-bounded, and proper function has a minimum [rockafellar2009variational, Theorem 1.9]. Assumption 4.5(ii) would often hold when the influence of $f$ is dominant compared to that of $g$. We will discuss this issue in more detail in Section A. Moreover, under Assumption 4.5(ii), Assumption 4.5(iii) always holds because $\Psi$ is bounded below from Assumption 3.1(iv) and the right-hand side of (3.6) is unbounded below. In this case, we have $\|t_{k}d^{k}\|\to 0$.
@@
where the last inequality holds because $\phi$ is $\sigma$-strongly convex. Let $\bar{x}\in\mathbb{R}^{n}$ be an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$ and let $\{x^{k_{j}}\}_{j\in\mathbb{N}}$ be a subsequence such that $x^{k_{j}}\to\bar{x}$ by the Bolzano–Weierstrass theorem. From (4.26) and the Cauchy–Schwarz inequality, we have
@@
If $\|\nabla f(x^{k})+\xi^{k}\|=0$, then $x^{k}$ becomes a stationary point. We assume $\|\nabla f(x^{k})+\xi^{k}\|>0$. By the triangle inequality, we have $\|\nabla f(x^{k})+\xi^{k}\|\leq\|\nabla f(x^{k})\|+\|\xi^{k}\|$. Since the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded, $\|\nabla f(x^{k})\|$ is bounded by the extreme value theorem and $\|\xi^{k}\|$ is also bounded (see, e.g.[Takahashi2022-ml, Theorem 1(ii)]). Thus, $\|\nabla f(x^{k})+\xi^{k}\|$ is bounded, i.e., due to (4.27), $d^{k}$ is also bounded. Thus, there exists a subsequence $\{d^{k_{j}}\}_{j\in\mathbb{N}}$ such that $d^{k_{j}}\to\bar{d}$ as $j\to\infty$ by the Bolzano–Weierstrass theorem. Then, by Lemma 4.6, the sequence $\{x^{k_{j}}+t_{k_{j}}d^{k_{j}}\}_{j\in\mathbb{N}}$ also converges to $\bar{x}$.
@@
which implies $d^{k_{j}}\to 0$. Because $f$ and $g$ are lower semicontinuous, from (4.24), we have
@@
where the first inequality holds due to the Cauchy–Schwarz inequality and $\nabla f$ being Lipschitz continuous on any compact subset from Assumption 4.8, and the second inequality holds due to (4.34). Moreover, using the inequality (4.26), we have
@@ 4.3 Global Convergence
where the second inequality holds because of the Lipschitz continuity of $\nabla f$ and $\nabla\phi$ on compact subsets from Assumption 4.8 and Assumption 3.1(i), and the last inequality holds from line 1 in Algorithm 1.
@@
    There exists $\bar{k}_{0}\in\mathbb{N}$ such that, for any $k\geq 1$, the following conditions hold:
@@
Suppose that Assumptions 3.1, 3.2, 4.1, 4.5, and 4.8 hold. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by ABPG-VMAW and let $\bar{x}$ be a stationary point of (1.1). Suppose further that $\Psi$ is a KL function with $\psi$ in the KL inequality (2.5) taking the form $\psi(s)=cs^{1-\theta}$ for some $\theta\in[0,1)$ and $c>0$. Then, the following statements hold:
@@ 5 Numerical Experiments
5.1 $\ell_{p}$-Regularized Least Squares Problem
We consider the sparse $\ell_{p}$-regularized least squares problem, where $p$ is slightly larger than 11[doi:10.1137/18M1194456, 7472557]:
$\min_{x\in\mathbb{R}^{n}}\quad\frac{1}{2}\|Ax-b\|^{2}+\frac{\theta_{p}}{p}\|x\|^{p}_{p},$ (5.1)
where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $\theta_{p}>0$. Let $g\equiv 0$. We also use $f$ and $\phi$ given by
@@ 5.1 $\ell_{p}$-Regularized Least Squares Problem
Note that $f$ and $\phi$ are $\mathcal{C}^{1}$ if $p>1$ while $\nabla f$ and $\nabla\phi$ are not globally Lipschitz continuous. Although we can choose any $\lambda>0$, we use $\lambda$ given by $\lambda<1/L$ if $(f,\phi)$ is $L$-smad (see, for more details, Remark 3.5). Note that our algorithm does not require the $L$-smad property.
@@
We compare ABPG-VMAW with ABPG [Takahashi], the proximal gradient algorithm (PG) with a constant step-size, and PG with line search (PGL). We set $c_{1}=0.99$, $c_{2}=0.999$, $\mu=0.9$, and $\eta=2$ for ABPG-VMAW and $c_{1}=0.99$ and $\delta=0.9$ for ABPG. Although $\nabla f$ is not Lipschitz continuous, PG uses the step-size $1/L$ given by (5.3). Note that PG does not guarantee global convergence. PGL searches $\lambda_{k}>0$ satisfying the descent lemma and uses the initial step-size $\lambda_{0}=1/L$ given by (5.3[doi:10.1137/1.9781611974997, p.283]. The initial point $x^{0}\in\mathbb{R}^{n}$ is generated from an i.i.d. normal distribution. The maximum number of iterations is 1000. The terminal condition is $\|x^{k}-x^{k-1}\|\leq 10^{-8}$.
@@
Next, we show the average performance of the four methods—ABPG-VMAW, ABPG, PG, and PGL—on the $\ell_{p}$-regularized least squares problem. Specifically, we selected combinations of $m$ and $n$ from the set $\{100,200\}\times\{1000,2000,5000\}$. For each combination, we generated 100 random instances: in each instance, we drew an $m\times n$ matrix $A$ and a ground truth vector $x^{*}\in\mathbb{R}^{n}$ with $10\%$ nonzero entries from an i.i.d. normal distribution. For each generated instance, we set $b=Ax^{*}$, $p=1.2$, and $\theta_{p}=0.1$. Table 1 presents the average performance, including the number of iterations, the accuracy of the recovered point, the objective values, and computation time, across 100 different instances. ABPG-VMAW outperformed ABPG, PG, and PGL. Moreover, ABPG-VMAW converged in fewer iterations and in a shorter amount of time than ABPG, PG, and PGL.
Figure 3 shows which of $y^{k}$ and $x^{k}+t_{k}d^{k}$ in line 1 of Algorithm 1 is selected in ABPG-VMAW. The red plot represents the values of $t_{k}$ that are actually adopted, showing that $x^{k}+t_{k}d^{k}$ is selected. Figure 3 indicates that the proposed method allows larger step-sizes, which may account for its convergence in fewer iterations than ABPG. Moreover, for ABPG-VMAW, we varied $\lambda$ by scaling $L$ by each value in $\{0.01,0.05,0.1,0.2,0.5,0.8,0.9,1.0,1.1,1.2,2.0,5.0,10.0,100\}$, where $L$ given by (5.3) and compared the results. The maximum number of iterations is 15000. The results are shown in Figure 4. When the value around $L$ is chosen, both the number of iterations and the computation time are reduced.
Figure 3: Trace of step-size parameter $t_{k}$ over iterations
Figure 4: Comparative analysis of algorithmic performance with varying $L$-smad parameter scaling
Table 1: Average number of iterations, objective function value, accuracy, and CPU time for ABPG-VMAW, ABPG, PG, and PGL using random instances of the $\ell_{p}$-regularized least squares problem (5.1) ($100$ instances for $m=100$, and $10$ instances for $m=1000$)
@@
100 1500 ABPG-VMAW 125 0.494 0.993 0.227
ABPG 847 0.494 0.993 1.022
PG 1000 6.218 1.068 0.624
PGL 1000 17.968 4.385 0.686
3000 ABPG-VMAW 219 0.503 0.997 1.623
ABPG 980 0.503 0.997 3.750
PG 1000 6.412 1.039 3.174
PGL 1000 43.325 7.374 3.285
5000 ABPG-VMAW 160 0.492 0.999 2.072
ABPG 1000 0.492 0.999 6.497
PG 1000 6.372 1.025 5.618
PGL 1000 79.301 10.345 5.799
1000 1500 ABPG-VMAW 108 0.496 1.000 0.397
ABPG 587 0.496 1.000 3.473
PG 1000 28.745 2.067 0.777
PGL 1000 9.874 2.326 1.037
3000 ABPG-VMAW 44 0.512 1.000 0.548
ABPG 504 0.512 1.000 7.251
PG 1000 36.533 1.958 3.775
PGL 1000 24.393 3.981 4.343
5000 ABPG-VMAW 47 0.498 1.000 1.025
ABPG 434 0.498 1.000 9.925
PG 1000 43.081 1.875 6.493
PGL 1000 45.076 5.704 7.414
@@ 5.2 Nonnegative Linear Inverse Problem
Given a nonnegative matrix $A\in\mathbb{R}^{m\times n}_{+}$ and a nonnegative vector $b\in\mathbb{R}^{m}_{+}$, the goal of nonnegative linear inverse problems is to recover a signal $x\in\mathbb{R}^{n}_{+}$ such that $Ax\simeq b$. Nonnegative linear inverse problems have been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de], as well as in optimization [Bauschke2017-hg, Takahashi]. To achieve the goal of nonnegative linear inverse problems, we focus on the convex optimization problem given by
$\min_{x\in\mathbb{R}^{n}_{+}}\quad D_{\rm{KL}}(Ax+b)+\theta_{1}\|x\|_{1},$ (5.4)
@@
Let $f(x)=D_{\rm{KL}}(Ax,b)$ and $g(x)=\theta_{1}\|x\|_{1}$. We use $\phi_{0}(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ as the kernel generating distance for BPG and $\phi_{1}(x)=\phi_{0}+\frac{1}{2}\|x\|^{2}$ as the kernel generating distance for our algorithm and ABPG. In this case, we also define $C=\operatorname{int}\operatorname{dom}\phi_{0}=\operatorname{int}\operatorname{dom}\phi_{1}=\mathbb{R}^{n}_{+}$. When $\sum_{i=1}^{m}a_{ij}=1$, the pair $(f,\phi_{0})$ is $1$-smad [Bauschke2017-hg] and the pair $(f,\phi_{1})$ is also $1$-smad [Takahashi]. We compare ABPG-VMAW with ABPG [Takahashi], PGL, and BPG. Those subproblems can be solved in closed form.
The problem setting is as follows. We generate the matrix $A\in\mathbb{R}^{m\times n}$ and the ground truth $x^{*}\in\mathbb{R}^{n}$, which has 5% nonzero elements, from an i.i.d. normal distribution. We set $b=Ax^{*}$. For $(n,m)=(200,500)$ and $\theta_{1}=0.05$, Figure 5 shows the objective function value $\Psi(x^{k})$ and the accuracy $\|x^{k}-x^{*}\|$ at each iteration on a logarithmic scale and Figure 6 shows those on the time axis.
@@
Figure 5: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (purple) on the nonnegative linear inverse Problem (5.4)
@@
Figure 6: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (purple) on the nonnegative linear inverse problem (5.4)
5.3 Phase Retrieval
We consider phase retrieval, i.e., recovering a signal $x\in\mathbb{R}^{n}$ such that $|\langle a_{i},x\rangle|^{2}\simeq b_{i}$ for $i=1,\ldots,m$, where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b_{i}\in\mathbb{R}$ is a observed magnitude. Phase retrieval has been studied in many applications, such as image processing [Candes2015-mk] and X-ray crystallography [Patterson1934-sg, Patterson1944-nm] as well as optimization [Bolte-2018, Takahashi2022-ml, Takahashi2025-py]. We address the following nonconvex optimization problem:
$\min_{x\in\mathbb{R}^{n}}\quad\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2}.$ (5.6)
Let $f(x)=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2}$. When we use $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$, $(f,\phi)$ is $L$-smad for any $L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)$ (see [Bolte-2018, Lemma 5.1]). We compare ABPG-VMAW with ABPG, PGL, and BPG. Those subproblems can be solved in closed form (see [Bolte-2018, Proposition 5.1] for the subproblem of BPG). We set $\lambda=1/L$, where $L=\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)$.
The problem setting is as follows. We generate $a_{i}\in\mathbb{R}^{n}$, $i=1,\ldots,m$, and the ground truth $x^{*}\in\mathbb{R}^{n}$ from an i.i.d. normal distribution. We set $b_{i}=|\langle a_{i},x^{*}\rangle|^{2}$. For $(n,m)=(200,1000)$, Figure 7 shows the objective function value $f(x^{k})$ and the accuracy $\|x^{k}-x^{*}\|$ at each iteration on a logarithmic scale, and Figure 8 shows those on the time axis.
Under this condition, ABPG-VMAW outperforms the other three methods in terms of the reduction in the objective function value per iteration and per unit time. ABPG-VMAW allows large step-sizes while ABPG and BPG use $\lambda=1/L$, which can be small, and PGL would estimate small step-sizes without Lipschitz continuous gradients. Moreover, the error with the ground truth is significantly smaller for ABPG-VMAW than for the other three methods.
(a) Objective function values
(b) Accuracy
Figure 7: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (purple) on the phase retrieval (5.6)
(a) Objective function values
(b) Accuracy
Figure 8: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (purple) on the phase retrieval (5.6)